OpenStudy (anonymous):

In a Geometric progression : a5 is 24 more than a3. if terms are Integer and +. a3+a5=?

hartnn (hartnn):

hi :) can you please post the exact question? " & +. a3+a5=?" is not making much sense...

OpenStudy (anonymous):

i mean if terms are integer and positive.a3+a5=????!

hartnn (hartnn):

so we have a5-a3 = 24 and we need a5+a3 right ? you know the formula for n'th term of arithmetic series ??

OpenStudy (anonymous):

right. this is a Geometric progression not arithmetic progression!

Parth (parthkohli):

Hi, I got it.

Parth (parthkohli):

Hartnn, continue. I'd type it later.

Parth (parthkohli):

Sorry lol. Wrong one.

hartnn (hartnn):

sorry i got disconnected. yeah, i mistyped it as arithmetic. we have ar^4-ar^2=24 and we need ar^4+ar^2 , hmmm... any more info given ??

OpenStudy (anonymous):

no :((

hartnn (hartnn):

i hope 1st or 2nd term was given, but let me try ar^2(r^2-1) = 24 since all terms are integer, r cannot be very large, trying r=2 , a *4 (3) = 24 a= 2 so, terms = 2,4,8,16,32,... a3 = 8, a5=32 difference = 32-8 = 24 as we needed!

hartnn (hartnn):

but this was guess and check method ...

hartnn (hartnn):

got what i did ?

hartnn (hartnn):

also r can't be 1, then, a= 0 r can't be 3, then a= fraction, not integer

Parth (parthkohli):

I had the same work :P

hartnn (hartnn):

i believe you parth :)

Parth (parthkohli):

Sarcasm... But something that was different is I had ar^2(r+1)(r-1) = 24 and I considered factors of 24 that were in the form r + 1 and r - 1 and then I eliminated 3,5 andc7 like you did

Parth (parthkohli):

Believe me now? :p

hartnn (hartnn):

yup, still do.